# Isometric rigidity of the Wasserstein space over the plane with the maximum metric

### v1

@@ 3.3. Diagonal rigidity of $\mathcal{W}_{1}(Q,d_{\mathrm{{m}}})$
is itself an isometry. Let us assume indirectly that $\Phi\circ\Psi_{\#}(\delta_{x_{3}})=\delta_{x_{4}}$ (which implies that and $\Phi\circ\Psi_{\#}(\delta_{x_{4}})=\delta_{x_{3}}$), and take a $\xi$ such that $\mathrm{supp}(\xi)\subseteq S_{3}$, $\xi\notin\{\delta_{x_{2}},\delta_{x_{3}}\}$. Since $d_{W_{p}}(\delta_{x_{1}},\xi)=1$, we have
$1=d_{W_{p}}(\delta_{x_{1}},\xi)=d_{W_{p}}(\Phi(\delta_{x_{1}}),\Phi(\xi))=d_{W_{p}}(\delta_{x_{1}},\Phi(\xi)),$
which implies that $\mathrm{supp}(\Phi(\xi))\subseteq S_{3}\cup S_{4}$. Similarly, $d_{W_{p}}(\delta_{x_{4}},\xi)=1$, and thus
$1=d_{W_{p}}(\delta_{x_{4}},\xi)=d_{W_{p}}(\Phi(\delta_{x_{4}}),\Phi(\xi))=d_{W_{p}}(\delta_{x_{3}},\Phi(\xi)),$
@@
(30) $d_{\mathrm{{m}}}((0,0),x)+d_{\mathrm{{m}}}(x,(1,1))\geq d_{\mathrm{{m}}}((0,0),(1,1))=1$
@@
The above inequality follows simply by the triangle applied for the metric $d_{\mathrm{{m}}}$. The characterization of the equality case is slightly more tricky. It is based on the fact that the line segment $L_{+}:t\to(t,t),t\in[0,1]$ is the only geodesic with respect to the metric $d_{\mathrm{{m}}}$ connecting the endpoints $(0,0)$ and $(1,1)$. This observation has the following consequence for measures: If $\mu\in\mathcal{W}_{1}(Q,d_{\mathrm{{m}}})$, then
(31) $d_{W_{1}}(\delta_{(0,0)},\mu)+d_{W_{1}}(\mu,\delta_{(1,1)})\geq 1,$
@@
(32) $\begin{split}d_{W_{1}}(\delta_{(0,0)},\mu)+d_{W_{1}}(\mu,\delta_{(1,1)})&=\int\limits_{Q}d_{\mathrm{{m}}}((0,0),x)~{}\mathrm{d}\mu(x)+\int\limits_{Q}d_{\mathrm{{m}}}(x,(1,1))~{}\mathrm{d}\mu(x)\\ &\geq\int\limits_{Q}~{}\mathrm{d}\mu(x)=1.\end{split}$
@@
$d_{\mathrm{{m}}}((0,0),x)+d_{\mathrm{{m}}}(x,(1,1))>1+\varepsilon,\ \textrm{for all}\ x\in B(x_{0},r).$
@@
$d_{W_{1}}(\delta_{(0,0)},\mu)+d_{W_{1}}(\mu,\delta_{(1,1)})=$
$=\int\limits_{\mathrm{supp}(\mu)}d_{\mathrm{{m}}}((0,0),x)\ ~{}\mathrm{d}\mu(x)+\int\limits_{\mathrm{supp}(\mu)}d_{\mathrm{{m}}}(x,(1,1))\ ~{}\mathrm{d}\mu(x)=$
$=\int\limits_{\mathrm{supp}(\mu)}[d_{\mathrm{{m}}}((0,0),x)+d_{\mathrm{{m}}}(x,(1,1))]\ ~{}\mathrm{d}\mu(x)=$
$=\int\limits_{\mathrm{supp}(\mu)\cap B(x_{0},r)}[d_{\mathrm{{m}}}((0,0),x)+d_{\mathrm{{m}}}(x,(1,1))]\ ~{}\mathrm{d}\mu(x)+$
$\hskip 42.67912pt+\int\limits_{\mathrm{supp}(\mu)\setminus B(x_{0},r)}[d_{\mathrm{{m}}}((0,0),x)+d_{\mathrm{{m}}}(x,(1,1))]\ ~{}\mathrm{d}\mu(x)$
$>(1+\varepsilon)\delta+(1-\delta)=1+\varepsilon\delta>1,$
@@
$d_{W_{1}}(\delta_{(0,0)},\Phi(\mu))+d_{W_{1}}(\Phi(\mu),\delta_{(1,1)})=d_{W_{1}}(\delta_{(0,0)},\mu))+d_{W_{1}}(\mu,\delta_{(1,1)}).$
@@
As mentioned at the beginning of the proof, the same argument shows that if we have $\mathrm{supp}(\mu)\subseteq L_{-}$, then it follows that $\mathrm{supp}\big{(}\Phi(\mu)\big{)}\subseteq L_{-}$ as well. Since $\mu_{0}=\delta_{(\frac{1}{2},\frac{1}{2})}$ has its support in $L_{+}\cap L_{-}$, and it is the unique measure with this property, we conclude, that $\Phi(\mu_{0})=\mu_{0}$. Now let us consider the Wasserstein space $\mathcal{W}_{1}(L_{+},d_{\mathrm{{m}}})$ of measures supported on $L_{+}$. By the above consideration, we have
@@
moreover, we know that $\Phi(\mu_{0})=\mu_{0}$ for the measure $\mu_{0}=\delta_{(\frac{1}{2},\frac{1}{2})}\in\mathcal{W}_{1}(L_{+},d_{\mathrm{{m}}})$. We cannot apply the characterization of Wasserstein isometries on a line segment (see [9, Theorem 2.5]) which implies that $\Phi(\mu)=\mu$ for all measures $\mu\in\mathcal{W}_{1}(L_{+},d_{\mathrm{{m}}})$. In the same way we can also conclude, that $\Phi(\mu)=\mu$ for all measures $\mu\in\mathcal{W}_{1}(L_{-},d_{\mathrm{{m}}})$.
@@ 3.4. Diagonal rigidity of $\mathcal{W}_{p}(Q,d_{\mathrm{{m}}})$ for $p>1$
In this final subsection, we consider the case $X=Q=[0,1]^{2}$ and $p>1$. In this case, we also have the statement:
@@
Let $\Phi$ be an isometry of the Wasserstein space $\mathcal{W}_{p}(Q,d_{\mathrm{{m}}})$. Since the Wasserstein space $\mathcal{W}_{p}([0,1],d_{|\cdot|})$ is isometrically rigid (see [9, Theorem 2.6]), Corollary 13 and the remark after that, implies that we can assume without loss of generality that $\Phi$ leaves every measure supported either on the right vertical side $[(1,0),(1,1)]$ or on the top horizontal side $[(0,1),(1,1)]$ fixed. We are going to show that $\Phi$ leaves every measure supported on the line diagonal segment $[(0,0),(1,1)]\subset Q$ fixed. The case of the other diagonal is analogous.
In the first step, we consider measures supported on the upper half of the diagonal, i.e. on the line segment $[(\frac{1}{2},\frac{1}{2}),(1,1)]$. We prove first that these measures will be fixed. In the second step we proceed to consider general measures supported on the full segment $[(0,0),(1,1)]$.
Let $\mu$ be a measure supported on the upper half of the diagonal of $Q$ (that is, on $[(\frac{1}{2},\frac{1}{2}),(1,1)]$), and consider the projections $p_{r}:\,(t,t)\mapsto(1,2t-1)$ and $p_{u}:(t,t)\mapsto(2t-1,1)$ that map $[(\frac{1}{2},\frac{1}{2}),(1,1)]$ onto the right vertical side $[(1,0),(1,1)]$ and the top horizontal side $[(0,1),(1,1)],$ respectively. Let us introduce the notation $\mu_{r}:=(p_{r})_{\#}(\mu)\quad\mbox{and}$ and $\mu_{u}:=(p_{u})_{\#}(\mu)$.
@@
Let $\nu\in\mathcal{W}_{p}(Q,d_{\mathrm{{m}}})$ be arbitrary, let $\pi_{(\mu_{r},\nu)}^{*}$ be an optimal transport plan between $\mu_{r}$ and $\nu$ with respect to the cost $c(x,y)=d_{\mathrm{{m}}}^{p}(x,y)$, let $\pi_{(\nu,\mu_{u})}^{*}$ be an optimal transport plan between $\nu$ and $\mu_{u}$ with respect to the same cost, and let $\pi_{(\mu_{r},\nu,\mu_{u})}^{*}\in\mathrm{Prob}(Q^{3})$ be the gluing of them — see the “gluing lemma” [15, Lemma 7.6] for further details of this construction. Moreover, set $\pi_{(\mu_{r},\mu_{u})}^{*}:=\left(\pi_{(\mu_{r},\nu,\mu_{u})}^{*}\right)_{13}.$ Now
@@
for all $(s,t)\in[\frac{1}{2},1]\times[\frac{1}{2},1].$ By the construction of $\mu_{r}$ and $\mu_{u},$ the couplings of these measures are in $1-1$ correspondence with the couplings of $\mu$ with itself — now we consider $\mu$ as a measure on $[\frac{1}{2},1].$ This correspondence is described as follows: if $\pi\in C(\mu_{r},\mu_{u}),$ then let us define $\tilde{\pi}\in C(\mu,\mu)$ by $\tilde{\pi}:=(p_{r}^{-1}\times p_{u}^{-1})_{\#}(\pi).$ Then
(39) $\begin{split}\iint_{Q^{2}}d_{\mathrm{{m}}}^{p}(x,z)~{}\mathrm{d}\pi(x,z)&=\iint_{[\frac{1}{2},1]^{2}}d_{\mathrm{{m}}}^{p}((1,2t-1),(2s-1,1))~{}\mathrm{d}\tilde{\pi}(t,s).\\ &\geq\iint_{[\frac{1}{2},1]^{2}}(1-(2t-1))^{p}~{}\mathrm{d}\tilde{\pi}(t,s)\\ &=\int\limits_{[\frac{1}{2},1]}(1-(2t-1))^{p}~{}\mathrm{d}\mu(t)\end{split}$
@@
$\iint_{Q^{2}}d_{\mathrm{{m}}}^{p}(x,z)~{}\mathrm{d}\pi(x,z)\geq\int\limits_{[\frac{1}{2},1]}(1-(2s-1))^{p}~{}\mathrm{d}\mu(s)$
which is saturated if and only if $t\geq s$ for $\tilde{\pi}-$a.e. $(s,t).$ Therefore, if $\pi$ is an optimal coupling of $\mu_{r}$ and $\mu_{u},$ then $\tilde{\pi}$ is supported on the diagonal of $[\frac{1}{2},1]^{2}$ which implies that $\tilde{\pi}=(id\times id)_{\#}(\mu)$. This means that $\pi=(p_{r}\times p_{u})_{\#}(\mu).$ So $(p_{r}\times p_{u})_{\#}(\mu)$ is the only optimal transport plan.
@@
$\mathrm{supp}(\pi_{(\mu_{r},\nu,\mu_{u})}^{*})\subset\left\{((1,2t-1),y,(2t-1,1))\,\middle|\,t\in[\frac{1}{2},1],\,y\in Q\right\}.$
@@
Now let us consider $\mu$, a probability measure supported on the main diagonal $[(0,0),(1,1)].$ The displacement interpolation given by
@@
is the unique geodesic line segment between $\mu_{0}=\delta_{(1,1)}$ and $\mu_{1}=\mu.$ Note that $\mu_{\frac{1}{2}}$ is supported on the “upper half of the diagonal” $[(\frac{1}{2},\frac{1}{2}),(1,1)]$ and hence preserved by $\Phi.$ Moreover, $\left(\mu_{s}\right)_{0\leq s\leq\frac{1}{2}}$ is the only geodesic line segment between $\delta_{(1,1)}$ and $\mu_{\frac{1}{2}},$ and the unique extension of this geodesic segment to the parameter domain $[0,1]$ is $\left(\mu_{s}\right)_{0\leq s\leq 1}.$ Therefore, the geodesic line segment $\left(\phi(\mu_{s})\right)_{0\leq s\leq 1}$ containing $\Phi(\mu_{0})=\Phi(\delta_{(1,1)})=\delta_{(1,1)}$ and $\Phi(\mu_{\frac{1}{2}})=\mu_{\frac{1}{2}}$ must coincide with $\left(\mu_{s}\right)_{0\leq s\leq 1},$ in particular, $\Phi(\mu)=\Phi(\mu_{1})=\mu_{1}=\mu.$

### v3

@@ 3.3. Diagonal rigidity of $\mathcal{W}_{1}(Q,d_{\mathrm{{m}}})$
is itself an isometry. Let us assume indirectly that $\Phi\circ\Psi_{\#}(\delta_{x_{3}})=\delta_{x_{4}}$ (which implies that and $\Phi\circ\Psi_{\#}(\delta_{x_{4}})=\delta_{x_{3}}$), and take a $\xi$ such that $\mathrm{supp}(\xi)\subseteq S_{3}$, $\xi\notin\{\delta_{x_{2}},\delta_{x_{3}}\}$. Since $d_{W_{p}}(\delta_{x_{1}},\xi)=2$, we have
$2=d_{W_{p}}(\delta_{x_{1}},\xi)=d_{W_{p}}(\Phi(\delta_{x_{1}}),\Phi(\xi))=d_{W_{p}}(\delta_{x_{1}},\Phi(\xi)),$
which implies that $\mathrm{supp}(\Phi(\xi))\subseteq S_{3}\cup S_{4}$. Similarly, $d_{W_{p}}(\delta_{x_{4}},\xi)=2$, and thus
$2=d_{W_{p}}(\delta_{x_{4}},\xi)=d_{W_{p}}(\Phi(\delta_{x_{4}}),\Phi(\xi))=d_{W_{p}}(\delta_{x_{3}},\Phi(\xi)),$
@@
(33) $d_{\mathrm{{m}}}((-1,-1),x)+d_{\mathrm{{m}}}(x,(1,1))\geq d_{\mathrm{{m}}}((-1,-1),(1,1))=2$
@@
The above inequality follows simply by the triangle inequality applied for the metric $d_{\mathrm{{m}}}$. The characterization of the equality case is slightly more tricky. It is based on the fact that the line segment $L_{+}:t\to(t,t),t\in[-1,1]$ is the only geodesic with respect to the metric $d_{\mathrm{{m}}}$ connecting the endpoints $(-1,-1)$ and $(1,1)$. This observation has the following consequence for measures: If $\mu\in\mathcal{W}_{1}(Q,d_{\mathrm{{m}}})$, then
(34) $d_{W_{1}}(\delta_{(-1,-1)},\mu)+d_{W_{1}}(\mu,\delta_{(1,1)})\geq 2,$
@@
(35) $\begin{split}d_{W_{1}}(\delta_{(-1,-1)},\mu)+d_{W_{1}}(\mu,\delta_{(1,1)})&=\int\limits_{Q}d_{\mathrm{{m}}}((-1,-1),x)~{}\mathrm{d}\mu(x)+\int\limits_{Q}d_{\mathrm{{m}}}(x,(1,1))~{}\mathrm{d}\mu(x)\\ &\geq\int\limits_{Q}2~{}\mathrm{d}\mu(x)=2.\end{split}$
@@
$d_{\mathrm{{m}}}((-1,-1),x)+d_{\mathrm{{m}}}(x,(1,1))>2+\varepsilon,\ \textrm{for all}\ x\in B(x_{0},r).$
@@
$d_{W_{1}}(\delta_{(-1,-1)},\mu)+d_{W_{1}}(\mu,\delta_{(1,1)})=$
$=\int\limits_{\mathrm{supp}(\mu)}d_{\mathrm{{m}}}((-1,-1),x)\ ~{}\mathrm{d}\mu(x)+\int\limits_{\mathrm{supp}(\mu)}d_{\mathrm{{m}}}(x,(1,1))\ ~{}\mathrm{d}\mu(x)=$
$=\int\limits_{\mathrm{supp}(\mu)}[d_{\mathrm{{m}}}((-1,-1),x)+d_{\mathrm{{m}}}(x,(1,1))]\ ~{}\mathrm{d}\mu(x)=$
$=\int\limits_{\mathrm{supp}(\mu)\cap B(x_{0},r)}[d_{\mathrm{{m}}}((-1,-1),x)+d_{\mathrm{{m}}}(x,(1,1))]\ ~{}\mathrm{d}\mu(x)+$
$\hskip 42.67912pt+\int\limits_{\mathrm{supp}(\mu)\setminus B(x_{0},r)}[d_{\mathrm{{m}}}((-1,-1),x)+d_{\mathrm{{m}}}(x,(1,1))]\ ~{}\mathrm{d}\mu(x)$
$>(2+\varepsilon)\delta+2(1-\delta)=2+\varepsilon\delta>2,$
@@
$d_{W_{1}}(\delta_{(-1,-1)},\Phi(\mu))+d_{W_{1}}(\Phi(\mu),\delta_{(1,1)})=d_{W_{1}}(\delta_{(-1,-1)},\mu))+d_{W_{1}}(\mu,\delta_{(1,1)}).$
@@
As mentioned at the beginning of the proof, the same argument shows that if we have $\mathrm{supp}(\mu)\subseteq L_{-}$, then it follows that $\mathrm{supp}\big{(}\Phi(\mu)\big{)}\subseteq L_{-}$ as well. Since $\mu_{0}=\delta_{(0,0)}$ has its support in $L_{+}\cap L_{-}$, and it is the unique measure with this property, we conclude, that $\Phi(\mu_{0})=\mu_{0}$. Now let us consider the Wasserstein space $\mathcal{W}_{1}(L_{+},d_{\mathrm{{m}}})$ of measures supported on $L_{+}$. By the above consideration, we have
@@
moreover, we know that $\Phi(\mu_{0})=\mu_{0}$ for the measure $\mu_{0}=\delta_{(0,0)}\in\mathcal{W}_{1}(L_{+},d_{\mathrm{{m}}})$. We cannot apply the characterization of Wasserstein isometries on a line segment (see [10, Theorem 2.5]) which implies that $\Phi(\mu)=\mu$ for all measures $\mu\in\mathcal{W}_{1}(L_{+},d_{\mathrm{{m}}})$. In the same way we can also conclude, that $\Phi(\mu)=\mu$ for all measures $\mu\in\mathcal{W}_{1}(L_{-},d_{\mathrm{{m}}})$.
@@ 3.4. Diagonal rigidity of $\mathcal{W}_{p}(Q,d_{\mathrm{{m}}})$ for $p>1$
In this final subsection, we consider the case $X=Q=[-1,1]^{2}$ and $p>1$. In this case, we also have the statement:
@@
Let $\Phi$ be an isometry of the Wasserstein space $\mathcal{W}_{p}(Q,d_{\mathrm{{m}}})$. Since the Wasserstein space $\mathcal{W}_{p}([-1,1],d_{|\cdot|})$ is isometrically rigid, Corollary 14 and the remark after that, implies that we can assume without loss of generality that $\Phi$ leaves every measure supported either on the right vertical side $[(1,-1),(1,1)]$ or on the top horizontal side $[(-1,1),(1,1)]$ fixed. We are going to show that $\Phi$ leaves every measure supported on the line diagonal segment $[(-1,-1),(1,1)]\subset Q$ fixed. The case of the other diagonal is analogous.
In the first step, we consider measures supported on the upper half of the diagonal, i.e. on the line segment $[(0,0),(1,1)]$. We prove first that these measures will be fixed. In the second step we proceed to consider general measures supported on the full segment $[(-1,-1),(1,1)]$.
Let $\mu$ be a measure supported on the upper half of the diagonal of $Q$ (that is, on $[(0,0),(1,1)]$), and consider the projections $p_{r}:\,(t,t)\mapsto(1,2t-1)$ and $p_{u}:(t,t)\mapsto(2t-1,1)$ that map $[(0,0),(1,1)]$ onto the right vertical side $[(1,-1),(1,1)]$ and the top horizontal side $[(-1,1),(1,1)],$ respectively. Let us introduce the notation $\mu_{r}:=(p_{r})_{\#}(\mu)\quad\mbox{and}$ and $\mu_{u}:=(p_{u})_{\#}(\mu)$.
@@
Let $\nu\in\mathcal{W}_{p}(Q,d_{\mathrm{{m}}})$ be arbitrary, let $\pi_{(\mu_{r},\nu)}^{*}$ be an optimal transport plan between $\mu_{r}$ and $\nu$ with respect to the cost $c(x,y)=d_{\mathrm{{m}}}^{p}(x,y)$, let $\pi_{(\nu,\mu_{u})}^{*}$ be an optimal transport plan between $\nu$ and $\mu_{u}$ with respect to the same cost, and let $\pi_{(\mu_{r},\nu,\mu_{u})}^{*}\in\mathrm{Prob}(Q^{3})$ be the gluing of them — see the “gluing lemma” [16, Lemma 7.6] for further details of this construction. Moreover, set $\pi_{(\mu_{r},\mu_{u})}^{*}:=\left(\pi_{(\mu_{r},\nu,\mu_{u})}^{*}\right)_{13}.$ Now
@@
for all $(s,t)\in[0,1]\times[0,1].$ By the construction of $\mu_{r}$ and $\mu_{u},$ the couplings of these measures are in $1-1$ correspondence with the couplings of $\mu$ with itself — now we consider $\mu$ as a measure on $[0,1].$ This correspondence is described as follows: if $\pi\in C(\mu_{r},\mu_{u}),$ then let us define $\tilde{\pi}\in C(\mu,\mu)$ by $\tilde{\pi}:=(p_{r}^{-1}\times p_{u}^{-1})_{\#}(\pi).$ Then
(42) $\begin{split}\iint_{Q^{2}}d_{\mathrm{{m}}}^{p}(x,z)~{}\mathrm{d}\pi(x,z)&=\iint_{[0,1]^{2}}d_{\mathrm{{m}}}^{p}((1,2t-1),(2s-1,1))~{}\mathrm{d}\tilde{\pi}(t,s).\\ &\geq\iint_{[0,1]^{2}}(1-(2t-1))^{p}~{}\mathrm{d}\tilde{\pi}(t,s)\\ &=\int\limits_{[0,1]}(1-(2t-1))^{p}~{}\mathrm{d}\mu(t)\end{split}$
@@
$\iint_{Q^{2}}d_{\mathrm{{m}}}^{p}(x,z)~{}\mathrm{d}\pi(x,z)\geq\int\limits_{[0,1]}(1-(2s-1))^{p}~{}\mathrm{d}\mu(s)$
which is saturated if and only if $t\geq s$ for $\tilde{\pi}-$a.e. $(s,t).$ Therefore, if $\pi$ is an optimal coupling of $\mu_{r}$ and $\mu_{u},$ then $\tilde{\pi}$ is supported on the diagonal of $[0,1]^{2}$ which implies that $\tilde{\pi}=(id\times id)_{\#}(\mu)$. This means that $\pi=(p_{r}\times p_{u})_{\#}(\mu).$ So $(p_{r}\times p_{u})_{\#}(\mu)$ is the only optimal transport plan.
@@
$\mathrm{supp}(\pi_{(\mu_{r},\nu,\mu_{u})}^{*})\subset\left\{((1,2t-1),y,(2t-1,1))\,\middle|\,t\in[0,1],\,y\in Q\right\}.$
@@
Now let us consider $\mu$, a probability measure supported on the main diagonal $[(-1,-1),(1,1)].$ The displacement interpolation given by
@@
is the unique geodesic line segment between $\mu_{0}=\delta_{(1,1)}$ and $\mu_{1}=\mu.$ Note that $\mu_{\frac{1}{2}}$ is supported on the “upper half of the diagonal” $[(0,0),(1,1)]$ and hence preserved by $\Phi.$ Moreover, $\left(\mu_{s}\right)_{0\leq s\leq\frac{1}{2}}$ is the only geodesic line segment between $\delta_{(1,1)}$ and $\mu_{\frac{1}{2}},$ and the unique extension of this geodesic segment to the parameter domain $[0,1]$ is $\left(\mu_{s}\right)_{0\leq s\leq 1}.$ Therefore, the geodesic line segment $\left(\phi(\mu_{s})\right)_{0\leq s\leq 1}$ containing $\Phi(\mu_{0})=\Phi(\delta_{(1,1)})=\delta_{(1,1)}$ and $\Phi(\mu_{\frac{1}{2}})=\mu_{\frac{1}{2}}$ must coincide with $\left(\mu_{s}\right)_{0\leq s\leq 1},$ in particular, $\Phi(\mu)=\Phi(\mu_{1})=\mu_{1}=\mu.$